# Handling Numeric Expressions in Automatic Speech Recognition

### v1

@@ 4.1 Limitations
The main limitation for the cascaded approach using an LLM is the ability of the LLM to follow the prompt correctly. This is expected to be handled even better for newer LLMs getting trained. For the fine-tuned ASR model the limitation is getting diverse data containing suitable numeric expression formatting.
@@
We also tried adapting the ASR model using batch weighting [16] and/or a factorization-based approach [2] together with the common voice training dataset. While the performance on the Common voice test sets improved, which is expected since the training and test datasets are more similar, the performance on the numeric expression formatting was slightly worse. Furthermore, we tried freezing the encoder or only adapting the final projection layer during the adaptation with the numeric expressions data. For both, the performance on the numeric expressions test data was slightly worse compared to not freezing any weights.
5 Conclusion
@@
In this paper, we tackled the problem of correctly formatting numeric expressions in ASR transcripts. Our experiments revealed that LLMs, particularly the latest models, deliver strong performance in recognizing and formatting numeric expressions. On the other hand, end-to-end models adapted with synthetic ASR data provide competitive performance.
6 Acknowledgements
This research was supported in part by a grant from Zoom Video Communications, Inc. The authors gratefully acknowledge the support.

### v2

@@ IV-A Limitations
The main limitation for the cascaded approach using an LLM (besides that the needed information has to be in the ASR hypothesis) is the ability of the LLM to follow the prompt correctly. This is expected to be handled even better for newer LLMs getting trained. For the fine-tuned ASR model, the limitation is to obtain diverse data containing suitable numeric expression formatting.
@@
We also tried to adapt the ASR model using batch weighting [20] and/or a factorization/LoRA-based approach [2, 21] together with the common voice training dataset. Although the performance on the Common voice test sets improved, which is expected since the training and test datasets are more similar, the performance on the numeric expression formatting was slightly worse. Furthermore, we tried freezing the encoder or only adapting the final projection layer during the adaptation with the numeric expressions data. For both, the performance on the numeric expression test data was slightly worse compared to not freezing any weights.
V Conclusion
@@
In this paper, we tackled the problem of correctly formatting numeric expressions in ASR transcripts. Our experiments revealed that LLMs, particularly the latest models, deliver strong performance in recognizing and formatting numeric expressions. However, end-to-end models adapted with synthetic ASR data provide competitive performance.
VI Acknowledgment
This research was supported in part by a grant from Zoom Video Communications, Inc. Furthermore, the projects on which this research is based were funded by the Federal Ministry of Education and Research (BMBF) of Germany under the number 01EF1803B (RELATER), the Horizon research and innovation program of the European Union under grant agreement No 101135798 (Meetween), and the KIT Campus Transfer GmbH (KCT) staff in accordance with the collaboration with Carnegie – AI. The authors gratefully acknowledge the support.